# Spectral Compressive Imaging via Chromaticity-Intensity Decomposition

### v1

@@ Abstract
In coded aperture snapshot spectral imaging (CASSI), the captured measurement entangles spatial and spectral information, posing a severely ill-posed inverse problem for hyperspectral images (HSIs) reconstruction. Moreover, the captured radiance inherently depends on scene illumination, making it difficult to recover the intrinsic spectral reflectance that remains invariant to lighting conditions. To address these challenges, we propose a chromaticity-intensity decomposition framework, which disentangles an HSI into a spatially smooth intensity map and a spectrally variant chromaticity cube. The chromaticity encodes lighting-invariant reflectance, enriched with high-frequency spatial details and local spectral sparsity. Building on this decomposition, we develop CIDNet—a Chromaticity-Intensity Decomposition unfolding network within a dual-camera CASSI system. CIDNet integrates a hybrid spatial-spectral Transformer tailored to reconstruct fine-grained and sparse spectral chromaticity and a degradation-aware, spatially-adaptive noise estimation module that captures anisotropic noise across iterative stages. Extensive experiments on both synthetic and real-world CASSI datasets demonstrate that our method achieves superior performance in both spectral and chromaticity fidelity. Code and models will be publicly available.
@@ 1 Introduction
Coded aperture snapshot spectral imaging (CASSI) has emerged as a promising architecture for capturing hyperspectral images (HSIs) in a single shot arce2013compressive ; wagadarikar2008single ; yuan2021snapshot . By jointly modulating the spectral cube with a coded aperture and dispersing it spatially through a prism, CASSI produces a 2D compressed measurement that encodes both spatial and spectral information. This compressive measurement fuses (shears) the spectral bands, making each pixel of the 2D sensor a mixture of many wavelengths. As a result, recovering the full 3D spectral image becomes a severely under-determined, ill-posed inverse problem.
@@
The difficulty comes mainly from two aspects. One is that spatial and spectral signals are highly overlapped and entangled in the compressed measurement. Many works attempt to address this through various priors or deep models, broadly categorized into four paradigms. Optimization-based methods liu2018rank ; yuan2016generalized introduce hand-crafted priors such as total variation or low-rank constraints. However, their performance is often limited in recovering spatial structures, especially under complex textures or noise. Plug-and-Play (PnP) approaches qiu2021effective ; zheng2021deep integrate powerful pre-trained denoisers into iterative solvers, yet these methods typically denoise each or few spectral band independently, neglecting spectral correlation and structure. Deep unfolding methods cai2022degradation ; dong2023residual ; li2023pixel ; meng2020gap ; zhang2024dual ; zhang2024improving bridge model-based and data-driven paradigms by learning iterative modules guided by the CASSI physics. End-to-end networks cai2022mask ; hu2022hdnet ; meng2020end ; miao2019net ; wang2025sˆ leverage CNNs or Transformers to directly infer the spectral cube from measurements. These deep learning-based frameworks implicitly exploit spatial-spectral dependencies and have shown promising performance. Recently, diffusion model pan2024diffsci ; wu2024latent ; zengspectral and Mambaqin2025detail have been used for spectral reconstruction. Nonetheless, all these approaches rely on network backbones to learn spatial-spectral features in an implicit manner. There lacks a clear and interpretable decomposition or quantitative structure that explicitly characterizes the physical roles of spatial and spectral components during reconstruction.
@@
The second challenge is that existing methods often overlook the impact of illumination. Since the captured spectral measurement is radiance-based, it inherently entangles the intrinsic surface reflectance with scene illumination. This coupling makes the reconstruction sensitive to lighting variations across time and environments, thereby limiting spectral accuracy. To address similar issues in the RGB image, prior works have explored intrinsic image decomposition bell2014intrinsic ; garces2022survey ; li2018cgintrinsics ; li2018learning and Retinex-based models wei2018deep ; wu2022uretinex to explicitly separate reflectance from illumination, enabling applications such as shadow removal and low-light enhancement. In the hyperspectral remote sensing community, several studies have also extended intrinsic decomposition to spectral reflectance and illumination separation jin2018intrinsic ; kang2014intrinsic ; yu2023hi , offering better invariance to lighting conditions. However, to the best of our knowledge, such decomposition has not yet been incorporated into CASSI reconstruction.
@@
where $\mathbf{C}$ denotes the chromaticity cube and $\mathbf{I}$ represents the intensity image. Notably, $\mathbf{C}$ exhibits several desirable properties: (i) spatially invariant to illumination, suppressing highlights and enhancing details in low-light regions; (ii) spectrally sparse with localized support (as illustrated latter); (iii) enriched with high-frequency texture, essential for fine-detail recovery. See Appendix C for more details. In contrast, the intensity component $\mathbf{I}$ captures the global illumination structure in the scene. It is interesting to note that the chromaticity exhibits more intrinsic characteristics of the sample compared to hyperspectral images. Hence, learning the chromaticity instead of HSIs seems to benefit the field more.
@@
    We develop an intensity-guided deep unfolding network that incorporates the chromaticity decomposition into unfolding algorithm. The network features a hybrid spatial-spectral Transformer architecture, where the encoder leverages window-based local spatial attention and the decoder employs sparse TopK spectral attention to capture localized spectral structures.
@@
    We introduce a degradation-aware, spatially-adaptive dual noise estimation module to model anisotropic noise across different reconstruction stages. This module enables each iteration to adaptively handle varying noise levels across spatial locations.
@@
Figure 2: (a) The architecture of our CIDNet with $K$ stages (iterations). (b) The CASSI system uses an intensity-guided mask to modulate the chromaticity. (c) Diagram of asymmetric backbone for our hybrid spatial-spectral transformer. (d) Sparse TopK spectral self-attention module.
@@ 2.1 Degradation Model of CASSI
intensity image can be approximated as a PAN image in dual-camera CASSI and the chromaticity is computed as the normalized spectral signature:
@@
where $\epsilon$ is a small constant to avoid division by zero. This decomposition separates the multiplicative effect of illumination $\mathbf{I}(u,v)$ from the spectral reflectance $\mathbf{C}(u,v,\lambda)$, which captures intrinsic scene properties. Importantly, $\mathbf{C}$ is invariant to changes in illumination intensity and direction (see Appendix A for derivation), enabling more robust modeling of reflectance and spectral reconstruction under varying lighting conditions. After decomposition, the CASSI measurement process can be modeled as follows. The hyperspectral cube $\mathbf{X}$ is modulated by a coded aperture $\mathbf{M}^{0}\in\mathbb{R}^{H\times W}$, resulting in a spatially coded cube:
$\mathbf{X}^{\prime}(u,v,\lambda)=\mathbf{C}(u,v,\lambda)\odot\mathbf{I}(u,v)\odot\mathbf{M}^{0}(u,v).$ (5)
Now we can treat $\mathbf{I}(u,v)\odot\mathbf{M}^{0}(u,v)$ as a new formation of coded mask $\mathbf{M^{\prime}}(u,v)=\mathbf{I}(u,v)\odot\mathbf{M}^{0}(u,v)$ incorporating the spatial intensity. Then following a typical CASSI formulation, the modulated cube $\mathbf{X}^{\prime}$ is then passed through a dispersive element that shifts each spectral band $\lambda_{n_{\lambda}}$ by a wavelength-dependent displacement $d(\lambda_{n_{\lambda}}-\lambda_{c})$ along the spatial axis (e.g., the $x$-axis). The sheared datacube can be expressed as:
@@
where $\mathbf{c}=\mathrm{vec}(\mathbf{C})$, $\mathbf{i}=\mathrm{vec}(\mathbf{I})$, $\bm{\Phi}$ is the sensing matrix determined by the modulation and dispersion process, and $\mathbf{n}$ is the vectorized noise term. Given that the intensity map $\mathbf{I}$ is known, we formulate the chromaticity-based measurement model as a standard linear inverse problem:
@@
where $\mathbf{c}=\mathrm{vec}(\mathbf{C})$ denotes the vectorized chromaticity, $\mathbf{H}$ is the effective sensing matrix that incorporates both the CASSI modulation-dispersion process and the known intensity modulation, and $\mathbf{n}$ is the vectorized noise.
@@ 2.2 Optimization Framework of CASSI
To characterize realistic imaging noise, we assume an anisotropic Gaussian noise model $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}=\mathrm{diag}(\sigma_{1}^{2},\ldots,\sigma_{M}^{2})$ represents the spatially varying noise covariance. Under a Bayesian framework, the posterior probability of the chromaticity $\mathbf{c}$ is given by $p(\mathbf{c}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{c})\cdot p(\mathbf{c})$, and the likelihood is:
@@
$\hat{\mathbf{c}}=\textstyle\arg\!\min_{\mathbf{c}}\left(-\frac{1}{2}(\mathbf{y}-\mathbf{H}\mathbf{c})^{\top}\bm{\Sigma}^{-1}(\mathbf{y}-\mathbf{H}\mathbf{c})\right)+\tau R(\mathbf{c}).$ (11)
When the noise is homoscedastic (i.e., $\bm{\Sigma}=\sigma^{2}\mathbf{I}$), the problem reduces to the common quadratic form: $\hat{\mathbf{c}}=\arg\!\min_{\mathbf{c}}\frac{1}{2}\|\mathbf{y}-\mathbf{H}\mathbf{c}\|_{2}^{2}+\tau R(\mathbf{c})$. We rewrite Eq. (11) using an auxiliary variable $\mathbf{z}$:
$\hat{\mathbf{c}},\hat{\mathbf{z}}=\arg\!\min_{\mathbf{c},\mathbf{z}}\left(-\frac{1}{2}(\mathbf{y}-\mathbf{H}\mathbf{c})^{\top}\bm{\Sigma}^{-1}(\mathbf{y}-\mathbf{H}\mathbf{c})\right)+\tau R(\mathbf{z}),\quad\text{s.t.}\quad\mathbf{c}=\mathbf{z}.$ (12)
Using the half-quadratic splitting (HQS) framework, Eq. (12) is minimized by solving the following subproblems iteratively:
$\mathbf{c}^{(k+1)}$ $=\textstyle\arg\!\min_{\mathbf{c}}\left(-\frac{1}{2}(\mathbf{y}-\mathbf{H}\mathbf{c})^{\top}\bm{\Sigma}^{-1}(\mathbf{y}-\mathbf{H}\mathbf{c})\right)+\frac{\mu}{2}\|\mathbf{c}-\mathbf{z}^{(k)}\|_{2}^{2},$ (13)
@@
where $\mu$ is a penalty parameter. The $\mathbf{c}$-subproblem in Eq. (13) is quadratic and has a closed-form solution:
$\mathbf{c}^{(k+1)}=\textstyle\left(\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{H}+\mu\mathbf{I}\right)^{-1}\left(\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{y}+\mu\mathbf{z}^{(k)}\right).$ (15)
In CASSI systems, $\mathbf{H}\mathbf{H}^{\top}$ is a diagonal matrix defined as $\mathbf{H}\mathbf{H}^{\top}\triangleq\mathrm{diag}\{h_{1},\ldots,h_{n}\}$. With $\bm{\Sigma}\triangleq\mathrm{diag}(\sigma_{1}^{2},\ldots,\sigma_{M}^{2})$ and detailed derivation (Appendix  B), we obtain a generalized form of gradient descent, which is expressed as,
@@
Define this gradient projection as $\mathbf{c}^{(k+1)}=\operatorname{proj}_{\bm{\Sigma}}(\cdot)$. Finally, we update $\mathbf{z}^{(k+1)}$ using any proximal operator depending on the prior $R(\cdot)$,
@@
The estimator $\mathcal{E}$ is implemented as a lightweight CNN that jointly captures spatial structure in $\mathbf{z}^{(k)}$ and the encoded degradation in $\mathbf{y}$. We denote $\bm{\Sigma}^{(k)}$ and $\omega^{(k)}=\tau^{(k)}/\mu^{(k)}$ as the noise map for gradient projection and proximal mapping (denoiser) respectively (Dual Noise-Estimation Module (DNEM), as we refered to). The estimated $\bm{\Sigma}^{(k)}$ reflects the anisotropic uncertainty in the current iterate, and modulates the linear update of $\mathbf{c}^{(k+1)}$ via Eq. (32), while $\omega^{(k)}$ controls the noise level fed into the proximal denoiser for $\mathbf{z}^{(k+1)}$. The final iterative process can be expressed as:
@@
Here, $\mu^{(k)}$ is omitted due to the usage of the network, $\mathbf{H}$ denotes the CASSI forward operator and $R(\cdot)$ is a regularization prior, which could be total variation, or a learned denoiser as in our experiments. This stage-adaptive formulation enables flexible and efficient recovery under spatially variant degradation patterns.
To validate the effectiveness of our chromaticity-intensity decomposition strategy, we explore two integration paradigms: a traditional model-based iterative scheme and a deep unfolding network. The classical iterative algorithm is described in the appendix, leveraging analytical priors and explicit update rules based on the degradation model. In contrast, our primary design adopts a learnable unfolding structure, as illustrated in Fig. 2. Each stage is composed of a learnable noise estimation, an analytical reconstruction step shown in Eq. (17) and a learned proximal denoiser. This framework offers the interpretability of traditional optimization while benefiting from the expressiveness and efficiency of deep networks.

### v2

@@ Abstract
In coded aperture snapshot spectral imaging (CASSI), the captured measurement entangles spatial and spectral information, posing a severely ill-posed inverse problem for hyperspectral images (HSIs) reconstruction. Moreover, the captured radiance inherently depends on scene illumination, making it difficult to recover the intrinsic spectral reflectance that remains invariant to lighting conditions. To address these challenges, we propose a chromaticity-intensity decomposition framework, which disentangles an HSI into a spatially smooth intensity map and a spectrally variant chromaticity cube. The chromaticity encodes lighting-invariant reflectance, enriched with high-frequency spatial details and local spectral sparsity. Building on this decomposition, we develop CIDNet—a Chromaticity-Intensity Decomposition unfolding network within a dual-camera CASSI system. CIDNet integrates a hybrid spatial-spectral Transformer tailored to reconstruct fine-grained and sparse spectral chromaticity and a degradation-aware, spatially-adaptive noise estimation module that captures anisotropic noise across iterative stages. Extensive experiments on both synthetic and real-world CASSI datasets demonstrate that our method achieves superior performance in both spectral and chromaticity fidelity. Code is released at: https://github.com/xiaodongwo/CIDNet.
@@ 1 Introduction
Coded aperture snapshot spectral imaging (CASSI) has emerged as a promising architecture for capturing hyperspectral images (HSIs) in a single shot Arce et al. (2013); Wagadarikar et al. (2008); Yuan et al. (2021). By jointly modulating the spectral cube with a coded aperture and dispersing it spatially through a prism, CASSI produces a 2D compressed measurement that encodes both spatial and spectral information. This compressive measurement fuses (shears) the spectral bands, making each pixel of the 2D sensor a mixture of many wavelengths. As a result, recovering the full 3D spectral image becomes a severely under-determined, ill-posed inverse problem.
@@
The difficulty comes mainly from two aspects. One is that spatial and spectral signals are highly overlapped and entangled in the compressed measurement. Many works attempt to address this through various priors or deep models, broadly categorized into four paradigms. Optimization-based methods Liu et al. (2018); Yuan (2016) introduce hand-crafted priors such as total variation or low-rank constraints. However, their performance is often limited in recovering spatial structures, especially under complex textures or noise. Plug-and-Play (PnP) approaches Qiu et al. (2021); Zheng et al. (2021) integrate powerful pre-trained denoisers into iterative solvers, yet these methods typically denoise each or few spectral band independently, neglecting spectral correlation and structure. Deep unfolding methods Cai et al. (2022b); Dong et al. (2023); Li et al. (2023); Meng et al. (2020a); Zhang et al. (2024a, b) bridge model-based and data-driven paradigms by learning iterative modules guided by the CASSI physics. End-to-end networks Cai et al. (2022a); Hu et al. (2022); Meng et al. (2020b); Miao et al. (2019); wang2025sˆ leverage CNN or Transformer to directly infer the spectral cube from measurements. These deep learning-based frameworks implicitly exploit spatial-spectral dependencies and have shown promising performance. Recently, diffusion model Pan et al. (2024); Wu et al. (2024); Zeng et al. and MambaQin et al. (2025) have been used for spectral reconstruction. Nonetheless, all these approaches rely on network backbones to learn spatial-spectral features in an implicit manner. There lacks a clear and interpretable decomposition or quantitative structure that explicitly characterizes the physical roles of spatial and spectral components during reconstruction.
@@
The second challenge is that existing methods often overlook the impact of illumination. Since the captured spectral measurement is radiance-based, it inherently entangles the intrinsic surface reflectance with scene illumination. This coupling makes the reconstruction sensitive to lighting variations across time and environments, thereby limiting spectral accuracy. To address similar issues in the RGB image, prior works have explored intrinsic image decomposition Bell et al. (2014); Garces et al. (2022); Li and Snavely (2018a, b) and Retinex-based models Wei et al. (2018); Wu et al. (2022) to explicitly separate reflectance from illumination, enabling applications such as shadow removal and low-light enhancement. In the hyperspectral remote sensing community, several studies have also extended intrinsic decomposition to spectral reflectance and illumination separation Jin et al. (2018); Kang et al. (2014); Yu et al. (2023), offering better invariance to lighting conditions. However, to the best of our knowledge, such decomposition has not yet been incorporated into CASSI reconstruction.
@@
where $\mathbf{C}$ denotes the chromaticity cube and $\mathbf{I}$ represents the intensity image. Notably, $\mathbf{C}$ exhibits several desirable properties: (i) spatially invariant to illumination, suppressing highlights and enhancing details in low-light regions; (ii) spectrally sparse with localized support (as illustrated latter); (iii) enriched with high-frequency texture, essential for fine-detail recovery. In contrast, the intensity component $\mathbf{I}$ captures the global illumination structure in the scene. It is interesting to note that the chromaticity exhibits more intrinsic characteristics of the sample compared to hyperspectral images. Hence, learning the chromaticity instead of HSIs seems to benefit the field more.
@@
    We develop an intensity-guided deep unfolding network that incorporates the chromaticity decomposition into unfolding algorithm. The network features a hybrid spatial-spectral Transformer (HSST) architecture, where the encoder leverages window-based local spatial attention (Spa-LWSA) and the decoder employs sparse TopK spectral attention (Spec-TKSA) to capture localized spectral structures.
@@
    We introduce a degradation-aware, spatially-adaptive dual noise estimation module (DNEM) to model anisotropic noise across different reconstruction stages. This module enables each iteration to adaptively handle varying noise levels across spatial locations.
@@
Figure 2: (a) The architecture of our CIDNet with $K$ stages (iterations). (b) The CASSI system uses an intensity-guided mask to modulate the chromaticity. (c) Diagram of asymmetric backbone for our hybrid spatial-spectral Transformer (HSST), with a local window spatial attention (Spa-LWSA) in Encoder and sparse TopK spectral attention module (Spec-TKSA) in Decoder. (d) Details of Spec-TKSA.
@@ 2.1 Degradation Model of CASSI
We found that intensity image can be approximated as a PAN image in dual-camera CASSI. A detailed proof in this PAN-Intensity Equivalence is provided in supplement materials. Hence, the chromaticity is computed as the normalized spectral signature:
@@
where $\epsilon$ is a small constant to avoid division by zero. This decomposition separates the multiplicative effect of illumination $\mathbf{I}(u,v)$ from the spectral reflectance $\mathbf{C}(u,v,\lambda)$, which captures intrinsic scene properties. Importantly, $\mathbf{C}$ is invariant to changes in illumination intensity and direction, enabling more robust modeling of reflectance and spectral reconstruction under varying lighting conditions (see supplement materials). After decomposition, the CASSI measurement process can be modeled as follows. The hyperspectral cube $\mathbf{X}$ is modulated by a coded aperture $\mathbf{M}\in\mathbb{R}^{H\times W}$, resulting in a spatially coded cube:
$\mathbf{X}^{\prime}(u,v,\lambda)=\mathbf{C}(u,v,\lambda)\odot\mathbf{I}(u,v)\odot\mathbf{M}(u,v).$ (5)
Now we can treat $\mathbf{I}(u,v)\odot\mathbf{M}{(u,v)}$ as a new formation of coded mask $\mathbf{M^{\prime}}(u,v)=\mathbf{I}(u,v)\odot\mathbf{M}(u,v)$ incorporating the spatial intensity, we call it intensity-guided mask. This leaves the chromaticity an unknown variable when the intensity is obtained beforehand. Following a typical CASSI formulation, the modulated cube $\mathbf{X}^{\prime}$ is then passed through a dispersive element that shifts each spectral band $\lambda_{n_{\lambda}}$ by a wavelength-dependent displacement $d(\lambda_{n_{\lambda}}-\lambda_{c})$ along the spatial axis (e.g., the $x$-axis). The sheared datacube can be expressed as:
@@
where $\mathbf{c}=\mathrm{vec}(\mathbf{C})$, $\mathbf{i}=\mathrm{vec}(\mathbf{I})$, $\bm{\Phi}$ is the sensing matrix determined by the modulation and dispersion process, and $\mathbf{n}$ is the vectorized noise term. Given that the intensity map $\mathbf{I}$ is known (PAN or RGB image in dual-camera CASSI scheme), we formulate the chromaticity-based measurement model as a standard linear inverse problem:
@@
where $\mathbf{c}$ denotes the vectorized chromaticity, $\mathbf{H}$ is the effective sensing matrix that incorporates both the CASSI modulation-dispersion process and the known intensity modulation, and $\mathbf{n}$ is the vectorized noise.
@@ 2.2 Optimization Framework of CASSI
Table 1: Data-consistency projection comparison.
Method Gradient projection updating
ISTA Zhang and Ghanem (2018) $\mathbf{c}^{k+1}=\mathbf{z}^{k}+\mathbf{H}^{\top}(\mathbf{y}-\mathbf{H}\mathbf{z}^{k})$
GAP Meng et al. (2020a) $\mathbf{c}^{k+1}=\mathbf{z}^{k}+\mathbf{H}^{\top}(\mathbf{H}\mathbf{H}^{\top})^{-1}(\mathbf{y}-\mathbf{H}\mathbf{z}^{k})$
HQS  Cai et al. (2022b) $\mathbf{c}^{k+1}=\mathbf{z}^{k}+\mathbf{H}^{\top}(\mathbf{H}\mathbf{H}^{\top}+\mu\mathbb{I})^{-1}(\mathbf{y}-\mathbf{H}\mathbf{z}^{k})$
Ours $\mathbf{c}^{k+1}=\mathbf{z}^{k}+\mathbf{H}^{\top}(\mathbf{H}\mathbf{H}^{\top}+\mu\bm{\Sigma})^{-1}(\mathbf{y}-\mathbf{H}\mathbf{z}^{k})$
To characterize realistic imaging noise, we assume an anisotropic Gaussian noise model $\mathbf{n}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma})$, where $\bm{\Sigma}=\mathrm{diag}(\sigma_{1}^{2},\ldots,\sigma_{n}^{2})$ is a diagonal covariance matrix whose diagonal entries $\sigma_{i}^{2}$ represent the noise variance at the $i$-th pixel. This implies that the noise is spatially varying but uncorrelated across pixels. Under a Bayesian framework, the posterior probability of the chromaticity $\mathbf{c}$ is given by $p(\mathbf{c}\mid\mathbf{y})\propto p(\mathbf{y}\mid\mathbf{c})\cdot p(\mathbf{c})$, and the likelihood is:
@@
$\hat{\mathbf{c}}=\textstyle\arg\!\min_{\mathbf{c}}\frac{1}{2}(\mathbf{y}-\mathbf{H}\mathbf{c})^{\top}\bm{\Sigma}^{-1}(\mathbf{y}-\mathbf{H}\mathbf{c})+\tau R(\mathbf{c}).$ (11)
When the noise is homoscedastic (i.e., $\sigma_{i}=1$), the problem reduces to the common quadratic form: $\hat{\mathbf{c}}=\arg\!\min_{\mathbf{c}}\frac{1}{2}\|\mathbf{y}-\mathbf{H}\mathbf{c}\|_{2}^{2}+\tau R(\mathbf{c})$. We rewrite Eq. (11) using an auxiliary variable $\mathbf{z}$:
$\hat{\mathbf{c}},\hat{\mathbf{z}}=\arg\!\min_{\mathbf{c},\mathbf{z}}\frac{1}{2}(\mathbf{y}-\mathbf{H}\mathbf{c})^{\top}\bm{\Sigma}^{-1}(\mathbf{y}-\mathbf{H}\mathbf{c})+\tau R(\mathbf{z}),\quad\text{s.t.}\quad\mathbf{c}=\mathbf{z}.$ (12)
Using the half-quadratic splitting (HQS) framework, Eq. (12) is minimized by solving the following data-consistency and data-prior subproblems iteratively:
$\mathbf{c}^{(k+1)}$ $=\textstyle\arg\!\min_{\mathbf{c}}\frac{1}{2}(\mathbf{y}-\mathbf{H}\mathbf{c})^{\top}\bm{\Sigma}^{-1}(\mathbf{y}-\mathbf{H}\mathbf{c})+\frac{\mu}{2}\|\mathbf{c}-\mathbf{z}^{(k)}\|_{2}^{2},$ (13)
@@
where $\mu$ is a penalty parameter and $k$ represent $k$th iteration. The $\mathbf{c}$-subproblem in Eq. (13) is quadratic and has a closed-form solution:
$\mathbf{c}^{(k+1)}=\textstyle\left(\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{H}+\mu\mathbb{I}\right)^{-1}\left(\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{y}+\mu\mathbf{z}^{(k)}\right),$ (15)
where $\mathbb{I}$ represent identity matrix. Note that $\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{H}$ is a fat matrix and $(\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{H}+\mu\mathbb{I})^{-1}$ will be difficult to compute and thus we simplify it based on the Sherman-Morrison-Woodbury formula,
$\left(\mathbf{H}^{\top}\bm{\Sigma}^{-1}\mathbf{H}+\mu\mathbb{I}\right)^{-1}=\mu^{-1}\mathbb{I}-\mu^{-2}\mathbf{H}^{\top}\left(\bm{\Sigma}+\mu^{-1}\mathbf{H}\mathbf{H}^{\top}\right)^{-1}\mathbf{H}.$ (16)
In CASSI systems, $\mathbf{H}\mathbf{H}^{\top}$ is a diagonal matrix defined as $\mathbf{H}\mathbf{H}^{\top}\triangleq\mathrm{diag}\{h_{1},\ldots,h_{n}\}$. With $\bm{\Sigma}\triangleq\mathrm{diag}(\sigma_{1}^{2},\ldots,\sigma_{n}^{2})$ and detailed derivation in supplement materials, we obtain a generalized form of gradient projection, which is expressed as,
@@
Define this gradient projection as $\mathbf{c}^{(k+1)}=\operatorname{proj}_{\bm{\Sigma}}(\cdot)$. Interestingly, we observed that this gradient projection resembles previous optimization-based methods but introduces a key change relates to the spatially-varying noise modeling. As summarized in Tab. 1, while traditional ISTA Zhang and Ghanem (2018), GAP Meng et al. (2020a), and HQS Cai et al. (2022b) methods employ static data-consistency steps with fixed regularization, our method proposes a dynamic, spatially-adaptive correction mechanism. Finally, we update $\mathbf{z}^{(k+1)}$ using any proximal operator depending on the prior $R(\cdot)$,
@@
The estimator $\mathcal{E}$ is implemented as a lightweight CNN that jointly captures spatial structure in $\mathbf{z}^{(k)}$ and the encoded degradation in $\mathbf{y}$. A detailed network module is found in supplement materials. We denote $\bm{\Sigma}^{(k)}$ and $\omega^{(k)}=\tau^{(k)}/\mu^{(k)}$ as the noise map for gradient projection and proximal mapping (denoiser) respectively (Dual Noise-Estimation Module (DNEM), as we refered to). The estimated $\bm{\Sigma}^{(k)}$ reflects the anisotropic uncertainty in the current iterate, and modulates the linear update of $\mathbf{c}^{(k+1)}$ via Eq. (17), while $\omega^{(k)}$ controls the noise level fed into the proximal denoiser for $\mathbf{z}^{(k+1)}$ via Eq. (18). The final iterative process can be expressed as:
@@
Here, $\mu^{(k)}$ is omitted due to the usage of the network, $R(\cdot)$ is a regularization prior, which could be total variation, or a learned denoiser as in our experiments. This stage-adaptive formulation enables flexible and efficient recovery under spatially variant degradation patterns.
To validate the effectiveness of our chromaticity-intensity decomposition strategy, we explore two integration paradigms: a traditional model-based iterative scheme and a deep unfolding network. The classical iterative algorithm is described in supplement materials, leveraging analytical priors and explicit update rules based on the degradation model. In contrast, our primary design adopts a learnable unfolding structure, as illustrated in Fig. 2. Each stage is composed of a learnable noise estimation, an analytical reconstruction step shown in Eq. (17) and a learned proximal denoiser. This framework offers the interpretability of traditional optimization while benefiting from the expressiveness and efficiency of deep networks.